sophie drew a triangle and reported the lengths is 3 inches, 10 inches and 9 inches can these three numbers represent the measures of the sides of a triangle? show how you determined the answer to the question?
Answer by jim_thompson5910(35256)


a = 3, b = 10, and c = 9 are the three sides of the triangle.

A triangle is only possible if we can add any two sides to have the sum be larger than the third side

a+b = 3+10 = 13 which is larger than c = 9
b+c = 10+9 = 19 which is larger than a = 3
a+c = 3+9 = 12 which is larger than b = 10
----------------------------------------------
So we've shown that the three inequalities below are true
a+b > c
b+c > a
a+c > b

Meaning that a triangle is possible with these three measurements.